OpenStudy (cedric):

Johnston High School has a total of 82 boys and girls who play sports. If the number of boys is 16 more than twice the number of girls, how many boys play sports at this high school?

OpenStudy (anonymous):

x-y=16 x+y=82 x :boy y:girl

OpenStudy (cedric):

so would i divide 82 by 2?

hero (hero):

That's wrong anyway. It's supposed to be x = boys y = girls x + y = 82 x = 2y + 16

OpenStudy (cedric):

so now do i divide the 16 by two??

hero (hero):

No, you do not divide 16 by anything. Subsitute 2y + 16 into the first equation, then solve for y. After solving for y, finding x should be easy.

OpenStudy (anonymous):

sr miss'' than twice"

OpenStudy (cedric):

y = 8 ??

hero (hero):

You could use a tutor.

OpenStudy (cedric):

):

OpenStudy (anonymous):

x= 82-y => second ep : 82-y =2y+16

OpenStudy (cedric):

66

OpenStudy (cedric):

oh sweet jesus its 42

OpenStudy (cedric):

@Hero

hero (hero):

x + y = 82 x = 2y + 16 Input since x = 2y + 16, replace x with 2y + 16 in the first equation to get (2y + 16) + y = 82 Remove parentheses: 2y + 16 + y = 82 Place like terms on the same side 2y + y = 82 - 16 Simplify: 3y = 66 y = 66/3 y = 22 Therefore x + 22 = 82 x = 82 - 22 x = 60 Therefore the number of boys is 60